OpenStudy (anonymous):

MEDAL AND FAN FOR BEST ANSWER!!! Solve. –6 ≥ 10 – 8x A.x ≤ 2 B.x ≥ –2 C.x ≥ 2 D.x ≤ –2

OpenStudy (anonymous):

All you do is solve it just like u would an equation..Do u know how to do that?

OpenStudy (anonymous):

go to wolfram alpha and paste the question into the box and it will automatically answer the question

OpenStudy (anonymous):

www.wolframalpha.com

OpenStudy (the_fizicx99):

.____. ^ –6 ≥ 10 – 8x Subtract both sides 10, \(\ \sf -6 \ge \cancel{10} - 8x\) \(\ \sf -10 \ge \cancel{-10}- 8x \) \(\ -16 \ge 8x \) Now isolate "x" by dividing both sides by 8: \(\ \sf \dfrac{-16}{-8} \ge \dfrac{\cancel{-8}x}{\cancel{-8}}\) When you divide or multiply by a negative number you switch the sign; \(\ \sf -2 \le x \)

OpenStudy (the_fizicx99):

Typo X_X -16/-8 = 2 not -2 D: